OpenStudy (anonymous):

Solve the linear equations 3x + 2y - 7 = 19 ------ (i) 4x - y + 27 = 4 --------(ii) 2x + 4y - 52 = 32 ------(iii) find x, y and z

OpenStudy (jack1):

where is z in the equations above...?

OpenStudy (jack1):

is it the -7, +27 and -52, or the 19, 4 and 32...?

OpenStudy (anonymous):

not sure also, that my problem also @jack1

OpenStudy (anonymous):

or use cramma's rule and let see what you will derive at

OpenStudy (jack1):

ok, assuming the 3rd number in every equation is a "z" value... you can solve it either as a matrix or as a series of simultaneous equations (they're both pretty much the same thing in the end, just different ways of approaching it, whatever you're more comfortable with really. anyway, i used matrix, and came up with: x= 372/115 (approx 3.24) y= 461/115 (approx 4) z= -21/115 (approx -0.18)

OpenStudy (anonymous):

thanks so much but there question is a sample question, here is what the got, am trying to use the sample to solve other questions. the sample question and working goes those:

OpenStudy (anonymous):

Example: Solve the linear equations 3x + 2y - 7 = 19 ------ (i) 4x - y + 27 = 4 --------(ii) 2x + 4y - 52 = 32 ------(iii) Solution: 3x + 2y - 7 = 19 --------- (i) Eq(i) x 4 -Eq(2) x 3 1 ly – l0z = 64 Eq(2) - Eq(3) x 2 -9y + 12z = -60 3x+2y-z= 19 1ly-l0z=64 Eq (4) x 9 - Eq (5) x 11 42 Z = -84 \ Z = -2 Substituting the value of Z in eq (4) you obtain lly+20=64 l ly + 20 = 64 11 y = 44 y = 4 Substituting the values of y and z in equation (i) You get 3x+8+2 = 19 3x = 9 x =3 \x =3,y =4 z =-2

OpenStudy (anonymous):

can you please teach me how you solve it, step by step explanation